# Attention as an Adaptive Filter

### v1

@@ 4.4 Implementation
Constants: stability $\epsilon$, causal mask $\mathbf{M}_{\text{causal}}$ Complex-valued projections:
@@
Compute normalized value directions: $\mathbf{U}_{\mathbf{v}}[:,i]=\mathbf{Z_{v}}[:,i]\big{/}\left(\|\mathbf{Z_{v}}[:,i]\|+\epsilon\right)$
@@
Tangential: $\mathbf{R}_{\text{tan}}^{2}=\|\mathbf{\tilde{V}}_{\mathbf{q}}\|_{2}^{2}\mathbf{1}_{m}^{\top}+\mathbf{1}_{m}\|\mathbf{\tilde{V}}_{\mathbf{k}}\|_{2}^{2}{}^{\top}-2\mathrm{Re}(\mathbf{E}_{\mathbf{qk}}\odot\mathbf{\tilde{V}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{V}}_{\mathbf{k}})$
@@
Geodesic step: $\mathbf{\bar{U}}_{\mathbf{v}}[:,i]=\left(\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\right)\Big{/}\left(\|\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\|+\epsilon\right)$
@@
$\mathbf{\bar{Z}_{v}}\leftarrow(1-\text{sigmoid}(\delta))\mathbf{Z_{v}}+\text{sigmoid}(\delta)\mathbf{\bar{Z}_{v}},\quad\mathbf{Z}_{\mathbf{v}}^{\text{pred}}=\mathbf{E_{v}}[:,1]\odot\mathbf{\bar{Z}}_{\mathbf{v}}$
@@
$\big{(}\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big{)}\leftarrow\mathcal{L}_{P}(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{\text{pred}}))$
@@ 4.5 Interpreting the Transformer as an MLE on a Hypersphere
where $\mathbf{\hat{u}}_{\mathbf{zs}ij}^{(k)}=\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}/\|\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\|=e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}$. Assuming that the transformed magnitude is approximately constant across time steps $e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\approx m_{\mathbf{zs}i}^{(k)}$, gives a simple precision-weighted average of $\mathbf{\hat{u}}$:
@@
$\mathbf{y}_{\mathbf{zs}i}^{(k)}\approx\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k)}}{\|\mathbf{y}_{\mathbf{zs}i}^{(k)}\|},$
@@
This structure — Norm, Attention, Add & Norm — closely resembles a Transformer.
@@
We present the following Transformer-inspired simplification of RT-AFA:
@@
Magnitude and unit direction:
$\mathbf{m}=\|\mathbf{Z}_{r}\|_{\text{rowwise}}\in\mathbb{R}^{1\times m}$
$\mathbf{U}_{r}=\gamma_{1}\mathbf{Z}_{r}\big{/}(\mathbf{m}+\epsilon)+\beta_{1}\quad(\text{element-wise division})$
Lift to complex:
$\mathbf{U}=[\mathbf{U}_{r};\mathbf{0}]\in\mathbb{C}^{2\times d\times m}$
Apply AFA attention:
$\mathbf{\hat{m}}=\mathrm{AFA}_{\mathrm{mag}}(\mathbf{m})$
$\mathbf{\hat{U}}=\mathrm{AFA}_{\mathrm{dir}}(\mathbf{U})$
@@
$\mathbf{m}_{y}=\|\mathbf{Y}\|_{\text{row-wise}}$
$\mathbf{\tilde{U}}=\gamma_{2}\mathbf{Y}\big{/}(\mathbf{m}_{y}+\epsilon)+\beta_{2}\quad(\text{element-wise division})$
@@ 5.1 Classical Filtering Methods
Norbert Wiener’s 1949 work on the Wiener Filter established the theoretical framework for minimum mean-square error estimation of signals from noisy observations, using second-order statistics (covariances) of the signal and noise (Wiener, 1964). Kalman introduced the discrete-time Kalman Filter, a recursive algorithm that optimally estimates the states of linear time-invariant stochastic systems from noisy measurements, achieving minimum mean-square estimation error (Kalman, 1960). Kalman & Bucy extended the approach to continuous-time systems with Gaussian noise (Kalman and Bucy, 1961). Rauch, Tung, and Striebel generalized the Kalman Filter to a batch smoothing framework, providing a maximum-likelihood estimator for the entire trajectory of a linear dynamical system by combining all measurements simultaneously, rather than recursively, which yields smoothed state estimates with minimum mean-square error over the full time horizon (Rauch et al., 1965). Extensions to nonlinear systems can be handled by the Extended Kalman Filter (EKF), which linearizes the dynamics and measurement functions about the current state estimate (Anderson and Moore, 1979). For introductions to stochastic processes and adaptive filtering, see (Jazwinski, 1970; Speyer, 2008; Haykin, 2002).
@@ 5.2 Robust and Adaptive Filtering
In many practical settings, the system and noise parameters required by the Kalman Filter are unknown or time-varying. Early work in adaptive signal processing began with Widrow & Hoff, who introduced the Least Mean Squares (LMS) algorithm, laying the foundation for recursive adaptation using the innovation signal (Widrow and Hoff, 1960). Schweppe provided early maximum likelihood foundations for estimating noise covariances in Gaussian signals, which became a cornerstone for statistically principled adaptive filtering (Schweppe, 1965). Mehra formalized adaptive Kalman Filtering by estimating unknown process and measurement noise covariances from the innovations sequence (Mehra, 1970), and compared innovation-based versus residual-based adaptation (Mehra, 1972). Iteratively reweighted least squares (IRLS) provides a framework for robust estimation by repeatedly solving weighted least-squares problems, with weights updated based on residuals; this approach was formalized in signal processing contexts by Burrus, Barreto, and Selesnick for FIR filter design (Burrus et al., 1994). Comprehensive treatments of stochastic estimation and adaptive methods include (Jazwinski, 1970) and (Maybeck and Siouris, 1980).
@@ 5.3 Neural Networks for Filtering Dynamical Systems
Goel and Bartlett investigate whether Transformers can represent classical Kalman Filters for linear time-invariant (LTI) systems. Their Transformer Filter uses softmax self-attention over nonlinear embeddings of past state estimates and observations. It effectively approximates the Kalman update indirectly, with the kernel weights mimicking the Kalman gain, and requires careful tuning of a temperature parameter to control approximation error (Goel and Bartlett, 2024). While this approach can produce state estimates arbitrarily close to the true Kalman Filter and approximate the behavior of an LQG controller, it does not naturally generalize to robust, time-varying, or non-Gaussian settings, nor does it explicitly propagate uncertainty in a structured way.
Several recent works integrate neural networks or attention into Kalman Filtering. KalmanNet uses a recurrent network to learn the Kalman gain for real-time filtering under partially known or nonlinear dynamics (Revach et al., 2022). The Neural Extended Kalman Filter parameterizes both process and observation models with neural networks for end-to-end latent state inference (Liu et al., 2023). The Attention Kalman Filter (AtKF) replaces the traditional Kalman gain with a self-attention network that learns to capture dependencies across past state and innovation sequences (Wang et al., 2024). A-KIT (2024) learns process noise covariance online to improve EKF predictions (Cohen and Klein, 2024). KalmanFormer integrates a Transformer network into the Kalman Filter framework to directly learn the Kalman gain from data. (Shen et al., 2025).
The present work is an extension of the MS Thesis, "Adaptive Filter Attention," which shows that a tensorized attention formulation exactly solves the MLE for a diagonalizable LTI system with Gaussian noise (Racioppo, 2025).
@@ 5.4 Neural State Space Models
State space models (SSMs) are a powerful framework for modeling sequences and time series data by representing system dynamics through latent states that evolve over time. A key challenge in sequence modeling is maintaining memory over long time horizons, and avoiding the vanishing and exploding gradient problem faced by recurrent neural networks.
@@
The HiPPO framework improves memory retention in SSMs by introducing recurrent memory with optimal polynomial projections, which allows the model to compress and retain important information from the entire history in a structured latent state space (Gu et al., 2020). The Structured State Space Sequence model (S4) designs efficient state space layers that scale linearly with sequence length, using structured representations of the system matrices (Gu et al., 2022).
@@
Subsequent works such as Mamba further enhance efficiency by introducing selective state space components that dynamically focus computation on relevant parts of the sequence (Gu and Dao, 2024). The S4ND model applies state space models to multidimensional signals such as images and videos, interpreting them as higher-dimensional spatiotemporal signals (Nguyen et al., 2022) . Goel et al. propose a multi-scale architecture based on S4 designed specifically for raw audio waveform generation, and address instability issues in SSM sequence modeling (Goel et al., 2022). Mamba Mixer combines structured state space models with selective token and channel mixing via learned gating, enabling efficient and adaptive modeling of long sequences by focusing on relevant elements and features (Behrouz et al., 2024). Finally, simplified variants of state space layers have been proposed to reduce model complexity and improve training stability while retaining the key benefits of state space modeling in sequence tasks (Smith et al., 2023).
@@
Both S4 and our model exploit a diagonal-plus-low-rank (DPLR) structure to reduce the cost of inverting a dense $N\times N$ matrix, using the Sherman-Morrison formula for efficiency: a diagonal part captures independent modes, and a rank-1 correction captures global interactions. In S4, the DPLR matrix models deterministic state dynamics, determining how past inputs influence future states. In our model, the DPLR structure appears in the noise covariance matrices in a stochastic differential equation (SDE) formulation, which allows us to represent how uncertainty propagates through the dynamics.
@@
Our formulation relies on complex-valued weight matrices. Prior work on complex-valued neural networks includes studies of complex-valued activation functions, parameter initialization strategies, and the use of Wirtinger calculus, which treats complex parameters as pairs of real variables to enable effective gradient computation (Trabelsi et al., 2018), (Barrachina et al., 2023),(Hammad, 2024).
@@ 5.5 Attention as a Structured State Space Computation
Dao and Gu establish that a state space model (SSM) transformation is mathematically equivalent to multiplication by an $N$-sequentially semiseparable (SSS) matrix. This equivalence means that the recurrence inherent in SSMs can be expressed as a structured matrix multiplication, which directly parallels the matrix form of self-attention in Transformers. In particular, the attention matrix in a Transformer can be interpreted as a low-rank or structured matrix operating along the sequence dimension, making it amenable to the same efficient algorithms developed for SSS matrices. (Dao and Gu, 2024). While this formulation offers computational benefits and is grounded in a linear state-space model, it does not explicitly propagate uncertainty through the sequence or perform adaptive, residual-based reweighting.

### v2

@@ 5.4 Algorithm
Constants: stability $\epsilon$, causal mask $\mathbf{M}_{\text{causal}}$
1exComplex-valued projections:
@@
Compute normalized value directions: $\mathbf{U}_{\mathbf{v}}[:,i]=\mathbf{Z_{v}}[:,i]\big/\left(\big|\mathbf{Z_{v}}[:,i]\big|+\epsilon\right)$
@@
Tangential: $\mathbf{R}_{\text{tan}}^{2}=\big|\mathbf{\tilde{V}}_{\mathbf{q}}\big|_{2}^{2}\mathbf{1}_{m}^{\top}+\mathbf{1}_{m}\big|\mathbf{\tilde{V}}_{\mathbf{k}}\big|_{2}^{2}{}^{\top}-2\mathrm{Re}(\mathbf{E}_{\mathbf{qk}}\odot\mathbf{\tilde{V}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{V}}_{\mathbf{k}})$
@@
Geodesic step: $\mathbf{\bar{U}}_{\mathbf{v}}[:,i]=\left(\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\right)\Big/\left(\big|\mathbf{Z}_{\mathbf{v}}[:,i]+\mathbf{\bar{V}}_{\mathbf{v}}[:,i]\big|+\epsilon\right)$
@@
$\mathbf{\bar{Z}_{v}}\leftarrow(1-\text{sigmoid}(\delta))\mathbf{Z_{v}}+\text{sigmoid}(\delta)\mathbf{\bar{Z}_{v}},\quad\mathbf{Z}_{\mathbf{v}}^{+}=\mathbf{E_{v}}[:,1]\odot\mathbf{\bar{Z}}_{\mathbf{v}}$
@@
$\big(\operatorname{Re}(\mathbf{Z_{p}}),\operatorname{Im}(\mathbf{Z_{p}})\big)\leftarrow\mathcal{L}_{P}(\operatorname{Re}(\mathbf{Z}_{\mathbf{v}}^{+}),\operatorname{Im}(\mathbf{Z}_{\mathbf{v}}^{+}))$
@@ 5.5 Interpreting the Transformer as an MLE on a Hypersphere
where $\mathbf{\hat{u}}_{\mathbf{zs}ij}^{(k)}=\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}/\big|\mathbf{\hat{z}}_{\mathbf{s}ij}^{(k)}\big|=e^{-\bm{\Lambda}_{\Omega}\Delta t_{ij}}\mathbf{u}_{\mathbf{zs}j}^{(k)}$. Assuming that the transformed magnitude is approximately constant across time steps $e^{-\alpha\Delta t_{ij}}m_{\mathbf{zs}j}^{(k)}\approx m_{\mathbf{zs}i}^{(k)}$, gives a simple precision-weighted average of $\mathbf{\hat{u}}$:
@@
$\mathbf{y}_{\mathbf{zs}i}^{(k)}\approx\mathbf{u}_{\mathbf{zs}i}^{(k)}+\mathbf{\bar{u}}_{\mathbf{zs}i}^{(k)},\qquad\mathbf{u}_{\mathbf{zs}i}^{(k+1)}=\frac{\mathbf{y}_{\mathbf{zs}i}^{(k)}}{\big|\mathbf{y}_{\mathbf{zs}i}^{(k)}\big|},$
@@
This structure — Norm, Attention, Add & Norm — closely resembles a Transformer, except that the normalization is applied in the eigenbasis (after the query/key/value projections), to complex-valued vectors.
@@
We propose the following Transformer-inspired simplification of RT-AFA:
@@
Complex projections:
$\big(\operatorname{Re}(\mathbf{Z_{q}}),\operatorname{Im}(\mathbf{Z_{q}})\big)\leftarrow\mathcal{L}_{Q}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big)$,
$\big(\operatorname{Re}(\mathbf{Z_{k}}),\operatorname{Im}(\mathbf{Z_{k}})\big)\leftarrow\mathcal{L}_{K}\big(\operatorname{Re}(\mathbf{Z}),\operatorname{Im}(\mathbf{Z})\big)$.
Magnitudes and unit directions (complex-valued):
$\mathbf{m}_{\mathbf{q}}=\big|\mathbf{Z}_{\mathbf{q}}\big|_{\text{rowwise}};\,\mathbf{m}_{\mathbf{k}}=\big|\mathbf{Z}_{\mathbf{k}}\big|_{\text{rowwise}};\,\mathbf{m}_{\mathbf{v}}=\big|\mathbf{Z}_{\mathbf{v}}\big|_{\text{rowwise}}\in\mathbb{R}^{1\times m}$
$\mathbf{U}_{\mathbf{q}}=\gamma_{1}\mathbf{Z}_{\mathbf{q}}\big/(\mathbf{m}_{\mathbf{q}}+\epsilon);\,\mathbf{U}_{\mathbf{k}}=\gamma_{1}\mathbf{Z}_{\mathbf{k}}\big/(\mathbf{m}_{\mathbf{k}}+\epsilon);\,\mathbf{U}_{\mathbf{v}}=\gamma_{1}\mathbf{Z}_{\mathbf{v}}\big/(\mathbf{m}_{\mathbf{v}}+\epsilon)$
Apply AFA attention (excluding projections):
$\mathbf{\hat{m}}=\mathrm{AFA}_{\mathrm{mag}}(\mathbf{m_{q}},\mathbf{m_{k}},\mathbf{m_{v}})$
$\mathbf{\hat{U}}=\mathrm{AFA}_{\mathrm{dir}}(\mathbf{U_{q}},\mathbf{U_{k}},\mathbf{U_{v}})$
@@
$\mathbf{m}_{y}=\big|\mathbf{Y}\big|_{\text{row-wise}}$
$\mathbf{\tilde{U}}=\gamma_{2}\mathbf{Y}\big/(\mathbf{m}_{y}+\epsilon)\quad(\text{element-wise division})$
@@ 6.1 Classical Filtering Methods
Norbert Wiener’s 1949 work on the Wiener Filter established the theoretical framework for minimum mean-square error estimation of signals from noisy observations, using second-order statistics (covariances) of the signal and noise (Wiener, 1964). Kalman introduced the discrete-time Kalman Filter, a recursive algorithm that optimally estimates the states of linear time-invariant stochastic systems from noisy measurements, achieving minimum mean-square estimation error (Kalman, 1960). Kalman & Bucy extended the approach to continuous-time systems with Gaussian noise (Kalman and Bucy, 1961). Rauch, Tung, and Striebel generalized the Kalman Filter to compute smoothed state estimates over the full trajectory of a linear dynamical system. Their method recursively combines forward-filtered estimates with backward-time updates to produce maximum-likelihood estimates of the states with minimum mean-square error over the entire time horizon (Rauch et al., 1965). Extensions to nonlinear systems can be handled by the Extended Kalman Filter (EKF), which linearizes the dynamics and measurement functions about the current state estimate (Anderson and Moore, 1979).
@@ 6.2 Robust and Adaptive Filtering
In many practical settings, the system and noise parameters required by the Kalman Filter are unknown or time-varying. Early work in adaptive signal processing began with Widrow & Hoff, who introduced the Least Mean Squares (LMS) algorithm, laying the foundation for recursive adaptation using the innovation signal (Widrow and Hoff, 1960). Schweppe provided early maximum likelihood foundations for estimating noise covariances in Gaussian signals (Schweppe, 1965). Mehra formalized adaptive Kalman Filtering by estimating unknown process and measurement noise covariances from the innovations sequence (Mehra, 1970). Iteratively reweighted least squares (IRLS) provides a framework for robust estimation by repeatedly solving weighted least-squares problems, with weights updated based on residuals; this approach was formalized in signal processing contexts by Burrus, Barreto, and Selesnick for FIR filter design (Burrus et al., 1994).
@@ 6.3 Neural Networks for Filtering Dynamical Systems
Goel and Bartlett investigate whether Transformers can represent classical Kalman Filters for linear time-invariant (LTI) systems. Their Transformer Filter uses softmax self-attention over nonlinear embeddings of past state estimates and observations. It effectively approximates the Kalman update indirectly, with the kernel weights mimicking the Kalman gain, and requires careful tuning of a temperature parameter to control approximation error (Goel and Bartlett, 2024).
Several recent works integrate neural networks or attention into Kalman Filtering (Shlezinger et al., 2025). KalmanNet uses a recurrent network to learn the Kalman gain for real-time filtering under partially known or nonlinear dynamics (Revach et al., 2022). The Neural Extended Kalman Filter parameterizes both process and observation models with neural networks for end-to-end latent state inference (Liu et al., 2023). The Attention Kalman Filter (AtKF) replaces the traditional Kalman gain with a self-attention network that learns to capture dependencies across past state and innovation sequences (Wang et al., 2024). A-KIT (2024) learns process noise covariance online to improve EKF predictions (Cohen and Klein, 2024). KalmanFormer integrates a Transformer network into the Kalman Filter framework to directly learn the Kalman gain from data. (Shen et al., 2025).
The present work extends the MS Thesis, "Adaptive Filter Attention," which demonstrates that a tensorized, residual- and precision-weighted attention formulation implements a maximum-likelihood estimator for a diagonalizable linear time-invariant SDE (Racioppo, 2025).
@@ 6.4 Neural State Space Models
State space models (SSMs) are a powerful framework for modeling sequences and time series data by representing system dynamics through latent states that evolve over time. A key challenge in sequence modeling is maintaining memory over long time horizons, and avoiding the vanishing and exploding gradient problem faced by recurrent neural networks (Somvanshi et al., 2025).
@@
The HiPPO framework improves memory retention in SSMs by introducing recurrent memory with optimal polynomial projections, which allows the model to compress and retain important information from the entire history in a structured latent state space (Gu et al., 2020). The Structured State Space Sequence model (S4) designs efficient state space layers that scale linearly with sequence length, using structured representations of the system matrices (Gu et al., 2022). The Toeplitz Neural Network (TNN) efficiently models sequences using relative position–encoded Toeplitz matrices and a lightweight Relative Position Encoder, enabling fast token mixing and extrapolation to much longer sequences with log-linear complexity (Qin et al., 2023).
@@
Subsequent works such as Mamba further improve efficiency by introducing selective state space components that dynamically focus computation on relevant parts of the sequence (Gu and Dao, 2024). Mamba Mixer combines structured state space models with selective token and channel mixing via learned gating, enabling efficient and adaptive modeling of long sequences by focusing on relevant elements and features (Behrouz et al., 2024).
@@
Both S4 and our model exploit a diagonal-plus-low-rank (DPLR) structure to reduce the cost of inverting a dense $d\times d$ matrix, using the Sherman-Morrison formula for efficiency: a diagonal part captures independent modes, and a rank-1 correction captures global interactions. In S4, the DPLR matrix models deterministic state dynamics, determining how past inputs influence future states. In our model, the DPLR structure appears in the noise covariance matrices in a stochastic differential equation (SDE) formulation, which allows us to represent how uncertainty propagates through the dynamics.
6.5 Complex-Valued Neural Networks
@@
Complex-valued neural networks (CVNNs) extend traditional real-valued ANNs by using complex numbers in weights, activations, and computations, naturally capturing rotations, phase relationships, and oscillatory dynamics that arise in linear algebra. Prior work includes studies of complex-valued activation functions, parameter initialization strategies, and the use of Wirtinger calculus, which treats complex parameters as pairs of real variables to enable effective gradient computation (Hirose, 2012; Bassey et al., 2021; Barrachina et al., 2023; Hammad, 2024).
Trabelsi et al. (2018) introduced practical components for training deep complex-valued neural networks—complex convolutions, batch normalization, weight initialization, and activation functions—and demonstrated their effectiveness on vision and audio tasks, showing competitive performance with real-valued networks (Trabelsi et al., 2018). Eilers and Jiang (2023) developed core components for a complex-valued transformer, including complex Scaled Dot-Product Attention and layer normalization, enabling fully complex-valued models that match real-valued performance while improving robustness to overfitting on complex-valued tasks (Eilers and Jiang, 2023). Ran-Milo et al. (2024) show that complex-valued SSMs can express oscillatory and long-range mappings with moderate dimensions and parameters, while real SSMs require much larger or exponentially scaled parameters for the same expressivity (Ran-Milo et al., 2024). Leng et al. develop a framework for complex-valued neural networks and transformers in wireless communications, showing that CVNNs require fewer layers than real-valued networks and introducing complex-valued embedding, encoding, decoding, and output modules (Leng et al., 2025).
@@ 6.6 Attention as a Structured State Space Computation
Dao and Gu establish a formal equivalence between the convolution kernel of a recurrent linear time-invariant (LTI) system and the attention weights in causal linear attention. They show that the state-space recurrence can be expressed as multiplication by an $N$-sequentially semiseparable (SSS) matrix, revealing that the attention matrix in a Transformer is effectively a low-rank, structured operator along the sequence dimension. This connection enables the use of fast algorithms developed for SSS matrices to accelerate attention. (Dao and Gu, 2024).
@@
6.7 Structuring Attention with Priors and Positional Biases
The original Transformer lacks built-in modeling of relative or absolute positions, relying instead on fixed additive sinusoidal positional encodings (Vaswani et al., 2023).
Many methods have proposed extending self-attention to efficiently incorporate relative position information. For example, Shaw et al. introduce a mechanism to encode relative distances between tokens directly into the attention computation (Shaw et al., 2018). Transformer-XL introduces relative positional encodings, enabling the model to generalize to sequences longer than those seen during training (Dai et al., 2019). Liu et al. treat the sequence of positional encodings as samples from a continuous trajectory governed by a Neural ODE. This formulation enforces smoothness, allows irregularly spaced positions via arbitrary sampling of time, and recovers sinusoidal encodings as a special case (Liu et al., 2020).
Su et al. introduce rotary positional encodings (RoPE), which encode relative position by applying a complex rotation to each query–key pair before computing their dot product. This makes attention scores depend only on relative displacement, yielding translation-equivariant attention (Su et al., 2023).
LeX Transformer extends RoPE with xPos, an exponential-decay stabilized rotary embedding, and blockwise masking, enabling Transformers to generalize smoothly to longer sequences without retraining (Sun et al., 2022). The Retentive Network (RetNet) combines each token’s hidden state with an exponentially decaying summary of previous tokens, using decay-masked attention to enforce causality and a gating mechanism to control past contributions and increase expressivity (Sun et al., 2023). Other models, including Dynamically Context-Sensitive Time-Decay Attention (Su et al., 2018) and RWKV also learn decay functions based on temporal proximity (Peng et al., 2023).
Some recent works focus on understanding the mathematical structure of rotary embeddings and extending them to higher dimensions or more flexible parameterizations. Yu et al. introduce ComRoPE, which replaces fixed 2D rotation matrices with trainable commuting angle matrices (Yu et al., 2025). Liu et al. use commuting skew-symmetric generators to define an $N$-dimensional RoPE, generalizing 1D and 2D forms and allowing learnable inter-dimensional interactions (Liu et al., 2025). CoPE (Amballa, 2025), and ComplexFormer (Shao et al., 2025) extend RoPE into the complex domain. Wu et al. observe that the same rotation trick used in Transformers (on queries and keys) can be applied to the matrices in an SSM formulation (Wu et al., 2025).
Press et al. introduce Attention with Linear Biases (ALiBi), which adds a distance-dependent bias to attention logits to enable length extrapolation (Press et al., 2022).
Several papers have explored element-wise (in the embedding dimension) attention mechanisms. Elliptical Attention replaces the traditional dot-product attention with a Mahalanobis distance-based mechanism (Nielsen et al., 2024). McCarter investigates replacing scaled dot-product attention with negative-log Euclidean distance (McCarter, 2023). Feng uses an exponentiated Euclidean distance as a similarity metric (Feng, 2025). Bacochina et al. replace matrix multiplications with learned spatial weight arrays applied directly via element-wise multiplication, focusing attention along rows and columns (Bacochina and de Souza, 2023). Wu et al. introduce the DA-Transformer, which reweights attention scores using learnable functions of token distances (Wu et al., 2021)